6x-4y=16
9x-6y=24
eliminate method

Answer by checkley79(3341)

6X-4Y=16 MULT. BY 3
9X-6Y=24 MULT. BY 2 & SUBTRACT.
18X-12Y=48
-18X+12Y=-48
------------------
0X+OY=0 NO SOUTION FOR X OR Y BECAUSE THE EQUATIONS ARE THE SAME.